Greenhab Report – February 5th

Crew 188 Sol 7 Green Hab Report 05FEB2018

Zac Trolley (Engineer), Guest writing for Julia De Marines (Crew Astronomer/GreenHab Officer)

05 February 2018

Environmental control: Heating

Ambient with door opening: Shade cloth on

Working hours in Green Hab: 11:00 – 11:30, 18:00 – 19:00

Outside temp at working hour: 17ºC

Inside temp at working hour: 25ºC

Inside temp H: 35ºC

Inside temp L: 14ºC

Inside humidity: 27 %

Inside humidity H: 79%

Inside humidity L: 18%

Hours of supplemental light: 05:00 – 9:59

Changes to the crops: The plants are following a predictable pattern of requiring sun and water, and producing green leaves. As the plants increase in size, we have been harvesting various parts of them to eat in our kitchen. On this 8th Sol, we dined on lettuce leaves and tasty herbs.

Daily water usage for crops: 48 litres

Time(s) of watering plants: 11:00 – 11:30 (16 litres), 18:00 – 19:00 (32 litres)

Changes to research plants: During an intensive research investigation it was discovered that one of the pillows in the VEGGIE OASYS experiment isn’t as wet as is desired. Further investigation is required in order to find the route cause of this discrepancy. The Executive Officer is working with the NASA PI in order to properly moisten the pillow.

Narrative: In order to better understand the needs and desires of the plants under our care, we have been closely monitoring their progress. We have been exploring alternative methods to increase plant yield, and have decided to ask the plants what they require to be at their maximum efficiency. To facilitate this, Zac Trolley and Sarah Jane Pell have undertaken an experimental procedure to breach the communication gap with the plants and create a dialogue with them directly. So far, our results are not promising.

Future needs and questions: We are awaiting the arrival of the tomato cages so we can replant the tall plants that require support to grow to their full potential. Please advise when they will arrive.

Sol Summary – February 5th

Crew 188 Sol 8 Summary Report 05FEB2018

Sol 8

Summary Title: Team Unity
Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer Mission Status: Nominal

Sol Activity Summary:
This morning our team awoke to a beautifully and thoughtfully arranged breakfast table with a special greeting written by two crewmembers who wanted to help our team start week 2 of our mission on a high note. Each mission brings with it a unique set of challenges and rewards and I am humbled and proud to be part of such a courageous and strong team.

Every morning at 08:30 we hold our Daily Planning Conference and one crewmember hosts a briefing of the upcoming EVA or team activity that will require the most coordination by all crew. We spent the mid-morning preparing for a 4-person EVA with the objective of scouting filming locations for a performing astronautics project by our Artist in Residence utilizing the Insta360 Pro camera. Two of our crew remained behind at the Hab to serve as HABCOM on the radio with the EVA team and keep up with science projects and tasks around the Hab.

Following the EVA we shared lunch and set about working on our individual projects. We had a productive afternoon followed by a period of relaxation. Our crew is currently in good spirits and communication, stress management and team work is at an all-time high. In the evenings when our crew is not too exhausted, we have started hosting a lecture series here on Mars. After our comms window closes and our crew finishes dinner and dishes, the next scheduled crewmember to present will spend about 30-45 minutes sharing a topic that they are either passionate about or teaching a useful skill. Our crew is comprised of six individuals with strong personal and professional backgrounds that make for informative evening sessions that bond and inspire us.

Ad Astra!
Renee Garifi
MDRS Crew 188 Executive Officer

Look Ahead Plan:
We will perform EVA #10 in the morning tomorrow and EVA #11 in the afternoon.

Anomalies in work:
None.

Weather:
Windy, but not like the high winds of yesterday. Skies were cloudy.

Crew Physical Status:
The crew is tired but in good spirits.

EVA:
Two EVAs planned for tomorrow.

Reports to file:
1. EVAs #10 and #11 request
2. EVA #9 Report
3. Ops Report
4. Sol Summary
5. Journalist Report
6. Green Hab Report
7. Daily Photos

Support Requested (acknowledged yesterday but listed as reminders): – Heater for science dome requested.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Water tank refills will be needed in the next few days.

Operations Report – February 5th

Crew 188 Operational Report 05/02/2018

Sol 8

Crew reporting: Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 84%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 5 Gallons

Water (trailer): 0 Gallons

Water (static): 150 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130788.5

Toilet tank emptied: Yes

ATVs Used: Yes, all four were used today.

Oil Added: None

ATV Fuel Used: 3 Gallons

# Hours the ATVs were Used today: 1

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:49

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Sub Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

EVA Report – February 5th

Crew 188 EVA #9 Summary Report 05FEB2018
Author’s name: Julia DeMarines, Crew Astronomer, Green Hab Operator

Purpose of EVA: Scouting of Crew 187’s recommended exploration area for chlorophyll sample collection and potential site for artistic Performing Astronautics and Bending Horizons projects using the Insta360 Pro Camera.

Location of EVA: “Despacito Rd” via 1101 off of Cow Dung Road UDM27 Coordinates: 519000 E, 4247500 N (the crew has mapped out several waypoints along potential path for confirmation)

Number of EVA Crew: 3
Participants: Ryan Kobrick (EVA CDR), Sarah Jane Pell, Julia DeMarines EVA Commander: Ryan Kobrick

Road(s) and routes per MDRS map:
1. Take East on Entrance Road (Driving)
2. Turn Right into Cow Dung Road (Driving)
3. Turn right on Crew 187’s “Despacito Rd” at 1101 marker (Driving) 4. Park at the end of the road near fence/gate
6. Resume walking towards West
7. Return by the same route

Mode of Travel: ATV and Pedestrian
Vehicles you will be using (if applicable): Blue ATVs 1, 2, 3, and Red ATV 4 Rovers used: None
Duration: 2 hours 41 minutes
EVA Departure Time: 10:26
EVA Return Time: 13:07

Summary:

We started up our ATV’s and drove staggered to our first waypoint at the intersection of Cow Dung road and “Despacito Rd” at the 1101 marker. Despacito Rd was much rockier than Cow Dung Rd and we advanced with caution around some high slope and eroded out areas. All four of us made it to our destination safely and parked the ATV’s. Next we walked around and explored the region for suitable spots for Dr. Sarah Jane Pell’s Performing Astronautics and Bending Horizons projects. Sarah Jane found a central spot to set up her Insta360º Pro camera and we recorded about 10 minutes of footage of the crew playing with bubbles at varying distances and heights around the camera. This made an excellent test run for Sarah Jane’s project, which she hopes to begin record tomorrow. Julia DeMarines collected several green samples for her Chlorophyll project.

Once the team was satisfied with their collections, footage, and exploration, we decided to head back home. This was not a difficult decision to make, as it was the hottest day on “Mars” since we arrived, and we were all feeling the heat. Not because we are all (obviously) outstandingly attractive but because the spherical helmet has a way of amplifying the Sun’s rays and we were wearing more than one layer of clothing. On the way back on Cow Dung Rd we stopped at a site of interest that we saw on the way out which looked like another excellent area for Sarah Jane’s Performing Astronautics and Bending Horizons project. We got off of our ATV’s and explored around for 10 minutes finding an even better location for filming. We marked a waypoint on Ryan’s GPS and headed back to the base. Once we had arrived we filled up the gas tanks on the ATV’s while we were waiting for Ryan’s dust project to be ready in the airlock. At 13:07 we entered the airlock, at 13:12 we entered the Hab, and at 13:31 we officially ended EVA #9 after removing our equipment, plugging them in, and cleaning our helmets. We will consider dressing lighter underneath our flight suits on future hot days as it was 20ºC by the time we returned.

Journalist Report – February 5th

Crew 188 Journalist Report 05FEB2018

Inspiring people using different language

SOL-8 Author’s Name: Tatsunari Tomiyama AHFP

[English]

The primary objective in this report is to educate and advertise our mission, crew 188, using different language(s). Our crews are mostly English speakers. I am the only crew member who can use an official space language other than English. By writing down English and Japanese, we are trying to educate other people who do not know English in Japan. As a secondary objective, the report is trying to inspire the Japan Mars Society. This report is the first time. Therefore, I would like to introduce our research projects while in the simulation. This report is written in English first for mission operation to help what I am writing and Japanese is followed.

Crew 188 is consisting by the International Space University (ISU) alumni. This is why Crew 188 is called Team ISU. I am also the alumni of Space Studies Program 2015 (SSP15). Other crew members also graduated with their masters and/or participated a program hosted by ISU previously. All of our member hold doctoral degree or a certain professional type of certification. Although I do not hold doctoral degree, I myself have Associate Human Factors Professional
certification by Board of Certification in Professional Ergonomics even though I am still studying for master’s degree.

Our crew member brought multiple research projects for the future space pioneering activity. The research is focusing on the Extra Vehicular Activity (EVA) and creating infrastructure at the station. Some of research is assigned by National Aeronautics and Space Administration (NASA) and significant aerospace research universities such as Embry-Riddle University. We brought over $30,000 for research equipments. In addition to research experiments, our team is trying to perform some public events. Before this report, super blue blood moon has been completed by Dr. Sarah Jane Pell.

We have spent about 1 week of missions while I am writing this report at the station. So far, we are doing well. Crew member statuses are good and there is no major issue on their health. All crew members spent time to complete their work or research and we are trying to be productive. In other words, we are very busy under this extreme environment at the Utah. My research has been accepted during this mission and collecting data on time. I do not know whether it is due to my research settings or not, our team situational awareness is relatively higher than other crew members in my opinion because everyone knows what we suppose to do, but we are recognizing other statuses at the same time.

[Japanese]

このレポートは、我々Crew188任務の宣伝ならびに、火星探査や宇宙関係に興味がある人に向けた事柄を英語以外で伝えることを主目的とする。Crew188は全員英語を何不自由なく使用することができるが、クルーの中で私だけが英語以外の言語を使うことが可能である。このレポートを通して、英語がわからない人でも我々Crew188の任務が理解できてもらえればと思っている。副次目的としては、日本語のレポートを通して日本火星協会をより応援することを期待している。今回は初めてのレポートであるため、クルーの構成と大まかな研究および実験の説明をする。サポートチームとの関係で、日本語でのレポートは最初に英語、そして日本語といった構成になる。

Crew188は国際宇宙大学(ISU)の卒業生で構成される。Crew188がチームISUと呼ばれる所以はこれから来る。私自身も実際にISUのスペーススタディプログラム2015の卒業生である。他のクルーメンバーもISUの大学院ないし、ISUが主催したプログラムに参加した経験を持つ。さらに、メンバー全員が博士号ないし、何かしらの専門家認定証を所持する。私自身はこれを書いている時点では大学院の生徒であるが、ヒューマンファクターズ(人的要因)を専門認定委員会から認定されている。

我々クルーメンバーは複数の研究、実験を持ち込んでおり、すべて将来の宇宙技術やそれに関する活動の創造するためのものである。そのため、Crew188の研究内容は主に船外活動(EVA)や火星探査を支える生活環境整備に関わる内容がほとんどを占める。いくらかの研究自体、米国航空宇宙局(NASA)やエンブルリドル航空大学(Embry-Riddle
Aeronautical University)のような航空宇宙を研究する重要な研究機関からの実験、研究依頼であり、実験及び研究するための備品費用でおおよそ$30,000ほど掛かっている。実験や研究活動以外にも教育及び宣伝活動を行っている。実際、先日の月食(スーパーブルーブラッドムーン)ではSarah Jane Pell教授が基地で月食を撮影し、その様子がインターネットに公表された。

このレポートを書いている時点でシミュレーションが始まって約1週間がたったところである。現時点での経過は総じて良好であると言える。各メンバー自体の体調や健康状態も良好である。メンバー全員がそれぞれの役割とデータ収集を行い、科学者として、ないし、一、プロフェッショナルとしての仕事をこなしている。言葉を言い換えれば、クルー全員がユタ州の人が生活していない場所で生活していながら非常に多忙であるといえる。しかしながら、個人の見解として、組織としての状況判断能力は他のメンバーと比べて高い判断している。というのも、たった数日ないし、1週間足らずの間でしか顔合わせをしていないのにも関わらず、メンバー全員がこなさなければならない事柄を行い、同時に他のメンバーの仕事や研究を認識しているからである。